OpenStudy (anonymous):

Evaluate the line integral \[\int_c F\cdot dr\] where c is the vector function \[\hat{r}(t)=t^3\hat i-t^2 \hat j+ t \hat k\] and \[F(x,y,z)=sinx\hat i + cosy \hat j+zx \hat k\] Here is how far I have gotten so far: \[\hat{r}'(t)=3t^2\hat i-2t \hat j+ \hat k\]

OpenStudy (lgbasallote):

you might want to use \cdot rather than \bullet hehe

OpenStudy (anonymous):

I'll change the Function F in terms of t's in just a second

OpenStudy (anonymous):

thanks @lgbasallote. Better? :P

OpenStudy (lgbasallote):

yes. yes it is.

OpenStudy (anonymous):

oh my, I did I similar problem 28 days ago http://openstudy.com/study#/updates/50060b97e4b06241806745b8 and I'm also looking at http://tutorial.math.lamar.edu/Classes/CalcIII/LineIntegralsVectorFields.aspx but I can't seem to remember... What do i substitute in for x and y, do I pick which one I want t to be?

OpenStudy (anonymous):

Hey i think I got it! \[x=t^3;y=-t^2;z=t\] YES?

OpenStudy (anonymous):

That would give me: \[F(r(t))=sin(t^3)\hat i +cos(-t^2)\hat j+t^4 \hat k \]

OpenStudy (anonymous):

\[\int_{t=0}^{t=1}\left(sin(t^3)\hat i +cos(-t^2)\hat j+t^4 \hat k\right)\cdot\left(3t^2\hat i -2t\hat j +\hat k \right) dt\]

OpenStudy (anonymous):

and now I would just do the dot product correct?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

\[\int_0^1(3t^2sin(t^3)-2tcos(-t^2)+t^4)dt\]

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

wait, I have to take the integral of several products?

OpenStudy (anonymous):

separate them\[\int_{0}^{1} 3t^2 \sin t^3 dt+...+...\]

OpenStudy (anonymous):

oh ok \[\int_0^1 3t^2sin(t^3)dt-\int_0^12tcos(-t^2)dt+\int_0^1t^4dt\] like this?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

oh integration by parts! duh =D

OpenStudy (anonymous):

\[-\cos t^3\] :-)

OpenStudy (turingtest):

no integration by parts you can do it all with u-subs

OpenStudy (turingtest):

do you get that?

OpenStudy (anonymous):

Oh I see it now!!!!

OpenStudy (turingtest):

cool :)

OpenStudy (anonymous):

u=t^2 du=2t dt for the second integral

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

u=t^3 and du=3t^2 dt for the first integral...LOL that took me a while!

OpenStudy (turingtest):

yeah, it's well set-up for the u-sub thing :)

OpenStudy (anonymous):

u cooked the problem

OpenStudy (anonymous):

\[-cos(1)+sin(1)+2\] My algebra is probably wrong...but Yeah @mukushla that was one long recipe!

OpenStudy (anonymous):

\[\large -\cos t^3]_{0}^{1}+\sin -t^2]_{0}^{1}+t^5/5]_{0}^{1}\]

OpenStudy (anonymous):

\[(-cos(1)+1)+(-sin(1)+1)+\frac 1 5\] \[-cos(1)-sin(1)+\frac{11}{5}\]

OpenStudy (anonymous):

oops sine of 0 is zero

OpenStudy (turingtest):

you have an extra +1 in there - -yep

OpenStudy (anonymous):

-cos(1)-sin(1)+ 6/5

OpenStudy (turingtest):

looks good to me :)

OpenStudy (anonymous):

sigh...finally! Thanks guys!

OpenStudy (turingtest):

welcome !

OpenStudy (anonymous):

:)